# and quantum information science

### v2

@@ 6.3.4 Random Matrix Theory
where $H$ is a Hamiltonian matrix acting on an $N$–dimensional Hilbert space, $V(H)=\sum_{n=0}^{\infty}v_{n}\mathrm{Tr}(H^{n})$ is a potential, and the partition sum $Z_{\beta_{D},N}$ normalizes the measure. [^26] In this section we return to using $N$ for the Hilbert space dimension rather than the number of degrees of freedom. The Dyson index is $\beta_{D}=1,2,4$ depending on whether we choose $H$ to be Hermitian, real and symmetric, or quaternionic, corresponding to unitary, orthogonal, or symplectic ensembles. If $V$ is quadratic, we say that the Random Matrix Theory (RMT) is described by the Gaussian Unitary, Orthogonal or Symplectic Ensembles (GUE,GOE, GSE). We are going to study time evolution and spread complexity of a generic state in the basis in which the Hamiltonian is drawn. Since the matrix is being drawn randomly, we can pick this to be the state $(1,0,0,\cdots)$ without loss of generality.
@@
[…]
@@
As explained in [23], we can use (104) to calculate the density of states from the potential of the RMT, and then we can use (71) from Sec. 6.1.4 to determine the coarse-grained Lanczos spectrum and the spread complexity. This construction is averaged over the ensemble, but as shown in [23], at large $N$ this gives the same result as coarse-graining over nearby Lanczos indices in a single draw from the ensemble, i.e., using the coarse-graining procedure outlined in Sec. 6.1.4. But we can do better than this and determine the distribution of Lanczos coefficients. For Gaussian RMTs, this distribution was worked out in [195]. The Jacobian arising from the basis transformation to get a tridiagonal Hamiltonian will be the same for any potential. The authors of [23] used this to show that for an RMT with an arbitrary potential like (102) the distribution of Lanczos coefficients is
@@
At large $N$ the distribution (105) is sharply peaked. So we can use the method of saddle points to extract the ensemble average (the saddle point) and the covariances (from the quadratic expansion around the saddle point). The ensemble average precisely reproduces (71), which we already derived through other means, and which applies to any theory for which we know the density of states. A formula for the covariances of the $a_{n}$ and $b_{n}$ is derived in [23, 24]. We will not repeat the derivation here for lack of space, but the result is
[…]
@@
[…]
@@
Among other things, these results show that the variance in $a(x)$ around the ensemble mean should be four times the variance in $b(x)$ for any RMT, where we are taking $x=n/N$ and coarse-graining as described in Sec. 6.1.4. The authors of [23] showed that these analytical formulae reproduce numerical results for RMTs with various potentials.
@@
Following [22], the spread complexity of the time-evolved thermofield-double state for a GUE Hamiltonian is plotted in Fig. 9 for various matrix sizes $N$ and inverse temperatures $\beta$. The dynamics exhibit four characteristic regimes: a linear ramp up to a peak that is exponential in the entropy (and hence linear in $N$, the Hilbert space dimension), followed by a slope down to a plateau. The authors of [22] showed that the peak and slope arise from eigenvalue correlations known as spectral rigidity in chaotic theories. These correlations can be removed by drawing eigenvalues at random from the Wigner semicircle distribution – this produces a spectrum without eigenvalue repulsion, like the Poisson spectra expected for integrable theories. The evolution with such a spectrum (Fig. 9 in light hues) shows only the ramp and plateau. Thus, the peak and slope are characteristic of late time chaos. The same patterns appear in the late time spread complexity of the chaotic SYK model [22].
@@
Figure 9: Quantum state complexity of the time-evolved thermofield-double state over an exponentially large period of time for different values of $N$ and $\beta$, as described in the main text. Dark Hues: GUE ensemble. Going from highest (blue) to lowest (yellow) curves we have $\beta=\{0,1,2,5,10\}$. In each case, we have plotted ensembles with $N=\{1024,1280,1536,1792,2048,2560,3072,3584,4096\}$. Complexity grows linearly to a peak, followed by a downward slope to a plateau. Light Hues: Ensemble with the same density of states as GUE, but without correlations between eigenvalues. In this case, the curves plateau without reaching a peak followed by a downward slope. Figure taken from [22].
@@
Quantum chaotic systems are conjectured to have spectra that are well-described by RMT statistics [57, 196]. In view of the results above, [24] conjectured that: Quantum chaotic systems display a Lanczos spectrum that is well described by RMT. As we discussed, the coarse-grained Lanczos spectrum can be determined from just the density of states. So the content of this conjecture lies in the correlations – it is saying that the covariances of the Lanczos coefficients are described by the RMT formulae stated above. To test this conjecture, we can numerically compute the Lanczos spectrum of integrable and chaotic spin chains such as
@@
where $X_{i}$, $Y_{i}$, $Z_{i}$ are Pauli spin operators on site $i$; $K$ is the number of spins; $A_{k}$, $B_{k}$ and $C_{k}$ are numerical parameters; and $H_{\text{site}}=0.5Z_{3}+0.3X_{3}$ is a one-site disorder operator that helps breaking parity and symmetry under the action of $\sum_{i}Z_{i}$. The model defined by $H_{\text{int1}}$ is integrable, while $H_{\text{cha1}}$ is chaotic as suggested by their level statistics, and have been explored in the literature [197, 175]. Indeed, as shown in [24], the analytical formulae developed above give an excellent description of the Lanczos coefficients in all cases, but the variances and covariances deviate significantly from the RMT formula for integrable spin chains, and in particular are much larger.
We can also write down an explicit formula for the height of the late time plateau in the spread complexity. First, let $\hat{n}=\sum_{n}n\lvert K_{n}\rangle\langle K_{n}\rvert$ be the position operator along the Krylov chain associated to an initial state $\lvert\phi\rangle$. Then, define the complexity of an energy eigenstate $\lvert E\rangle$ relative to $\lvert\phi\rangle$ as $\mathcal{C}_{E}=\langle E\rvert\hat{n}\lvert E\rangle$. A short calculation in [24] shows that the height of the late time plateau in the spread complexity will be
@@
where $p(E)=|\langle E|\phi\rangle|^{2}$ is the initial probability of eigenstate $\lvert E\rangle$. If $p(E)$ is continuous, we can replace $\mathcal{C}_{E}$ by the local average in a small window over energies to get a coarse-grained eigenstate complexity $\overline{\mathcal{C}_{E}}$. A calculation in [24] shows that $\overline{\mathcal{C}_{E}}$ is entirely determined by Lanczos coefficients, which are in turn determined in terms of the density of states following Eq. (71). In terms of the coarse grained eigenstate complexities, and in the large system limit, we can take a continuum limit of (109), which yields
[…]
This means that for large systems and initial states with continuous support in the energy basis, the plateau of the spread complexity will not discriminate between chaotic and integrable theories with the same average density of states. However, if $p(E)$ is “noisy” rather than smooth, and the noise is negatively correlated with the noise in $\mathcal{C}_{E}$, the plateau can be significantly lower than the averaged value in (110). This lowering can be larger for integrable theories because, as we discussed above, the variance in the Lanczos is higher, leading to a form of Anderson localization of energy eigenstates along the Krylov chain (see also [175]). So if an eigenstate is “centered” further from the beginning of the chain it tends to have less overlap with the initial state for an integrable theory. It would be interesting to investigate further how tuning the initial state leads to differences in the late time behavior for integrable versus chaotic theories because of variances and covariances in the Lanczos spectrum
These analyses have been extended to a broad class of integrable and chaotic models, including the SYK model and spin chains. Key aspects such as the presence or absence of the peak and slope [198], the peak height [199], and the system’s approach to the plateau [200] have been advocated as indicators of chaos. However, it seems that Krylov complexity [201] and spread complexity [202] are not sensitive to saddle-dominated scrambling (exponential growth due to unstable saddle points in phase space), at least for the initial states/operators studied in these papers, a phenomenon that OTOCs also fail to distinguish from genuine quantum chaos [203].
@@ 6.4 Open problems and directions
These quantities are known exactly in some QFTs. For example, in 2D CFTs, thermal two-point correlators are universally determined by Ward identities and are the same for both integrable and chaotic CFTs. In this case, there is an infinite sequence of $b_{n}$ Lanczos coefficients that grow linearly for large $n$ [204]. At least at first sight, this serves as a counterexample to the universal operator growth hypothesis. This issue was explored in [205, 206], where the authors argued that to meaningfully characterize operator growth in QFTs, one must first introduce appropriate cut-offs. They analyzed the effects of both UV and IR cut-offs, and concluded that the chaotic or scrambling properties of QFTs may be captured by $b_{n}$’s larger than the UV cut-off $\Lambda$. Meanwhile, the IR cut-off (e.g. modeled by introducing a mass) induced a separation (staggering) between even and odd Lanczos coefficients that grew linearly with different intercepts [206].
Thus, distinguishing characteristic features of chaotic and integrable operator growth in continuum QFTs remains an open problem. A promising approach could be to employ von Neumann algebras and algebraic QFTs. Some progress in this direction has been made in [207, 208], but a direct connection with Krylov-basis methods has not been established. Alternatively, we could compute the covariances of the Lanczos coefficients in QFT, in view of the conjecture in [23, 24] that their structure separates chaos from integrability. In theories with a discrete spectrum, it may also help to project onto a microcanonical band of energies (see also [209]) thereby rendering the accessible Hilbert space finite dimensional, thus enabling us to directly adopt the ideas and methods that have been developed for the latter.
@@
The relation between Krylov and Nielsen’s approaches has been debated in several works. In [210], the authors argued that Krylov complexity can be interpreted geometrically as a distance in a certain projection of the Fubini-Study geometry that is natural in quantum optics, making it amenable to a circuit complexity interpretation. However, as we already saw from the symmetry-based examples, Krylov complexity, as a function of time, is proportional to the volume in the Fubini-Study geometry. As a volume, rather than a length, it does not respect some properties of geodesics, such as the triangle inequality. Such properties were suggested in [211] as an obstruction to Krylov complexity serving as a distance measure. However, it is not clear that the Fubini-Study setting is the one in which we should seek a relation with Nielsen’s complexity. Indeed, [212] argued that, in certain setups, there is a relation between the time average of spread complexity of state evolution and an upper bound on Nielsen’s complexity with specific fine-tuned penalty factors. Recalling that the Nielsen approach has ambiguities in the choice of basis of generators and cost function, it remains possible that some judicious choice could establish a direct relation between the Nielsen and Krylov/spread approaches to complexity, at least for some specific choices of states. Perhaps useful progress could be achieved by formulating Krylov/spread complexity approach to general, time-dependent Hamiltonians (see progress in [213]). This would be closer in spirit to the starting point in Nielsen’s definitions and could allow us to use bounds on circuit complexity to constrain Krylov complexity. Indeed, if we treat the evolution parameter as a circuit time, we can mathematically think about the complexity of formation of states in the language of spread complexity [214]. Thus it remains an interesting problem to systematically relate the complexity of state/operator spreading to circuit complexity.
@@
It would be interesting to extend the spread/Krylov complexity approaches to new systems. First, the generalization to time-dependent Hamiltonians is important to achieve for the reasons described above, and because we would like to understand systems that are driven by external sources. Second, if we contemplate external sources, we should also consider open quantum systems, whose dynamics are controlled by Lindbladian evolution of the density matrix [215]. There is a substantial literature on the Linbladian formalism which is partly reviewed in [3]. The interaction of an external bath with an open system certainly affects the structure of entanglement spreading (see, e.g., [216]), and has generally been a subject of significant recent interest in condensed matter physics through the study of measurement-driven phase transitions (see, e.g., [217]). The spread complexity of measurement-induced non-unitary dynamics has been explored in [160]. Meanwhile, [218] have uncovered various bounds and relations pertaining to the state/operator complexity of purifying a mixed state. Another interesting avenue is to study the spread complexity of Carrollian theories of relevance for flat space holography. Likewise it would be interesting to study the kind of chiral theories that appear in the AdS/CFT dual to highly rotating extremal black holes [219, 220, 221] and the matrix models that appear as the dual to the 11-dimensional M-theory [222, 223, 224]. In this direction, spread complexity has been studied in the planar limit of holographic theories in [225].
@@
In Sec. 7.5 we will discuss how spread complexity of the double scaled SYK model is directly related to the length of the dual wormholes in two-dimensional Jackiw-Teitelboim gravity. But we could ask a broader question. Many lines of argument suggest that the microstates of black holes are immensely complex and may appear to be hidden behind a horizon simply because simple probes cannot discriminate between them – see, e.g., [226, 227, 228, 229]. Similarly, a number of authors have argued that in the AdS/CFT correspondence excitations behind the horizon are encoded in the CFT by exponentially complex operators [230, 231] in a non-isometric manner [232]. The information stored in Hawking radiation is also supposed to be encoded in an extremely complex manner, with the complexity expressed in the dual gravitational theory in terms of geometric obstructions preventing asymptotic observers from accessing the information [233]. At least in toy models of black hole evaporation like [234], such considerations can be explicitly realized and characterized in information-theoretic and complexity-theoretic terms [235]. But black holes can form from collapse of simple states like an ordered cloud of particles. So, the immense complexity of black hole microstates, and of the subsequent encoding of the initial state in the Hawking radiation, must be generated dynamically by evolution under the Hamiltonian of the theory. An important question is to understand how this happens. Perhaps the techniques of Krylov/spread complexity, or of Nielsen complexity from the previous section, can help. One recent result shows that if a holographic CFT time evolution operator is approximately pseudorandom or Haar random on a low energy subspace of states, there must be an event horizon in the dual theory of gravity [236]. Likewise, it seems that any region of AdS with sufficiently high holographic encoding complexity will be robustly protect from low-rank measurements, and hence be inaccessible to low-complexity observers [235]. It would be interesting to understand how the dynamical evolution of simple states can produce complex states with such properties.

### v3

@@ 6.3.4 Random Matrix Theory
where $H$ is a Hamiltonian matrix acting on an NN–dimensional Hilbert space, $V(H)=\sum_{n=0}^{\infty}v_{n}\mathrm{Tr}(H^{n})$ is a potential, and the partition sum $Z_{\beta_{D},N}$ normalizes the measure. [^29] In this section we return to using $N$ for the Hilbert space dimension rather than the number of degrees of freedom. The Dyson index is $\beta_{D}=1,2,4$ depending on whether we choose $H$ to be Hermitian, real and symmetric, or quaternionic, corresponding to unitary, orthogonal, or symplectic ensembles. If $V$ is quadratic, we say that the Random Matrix Theory (RMT) is described by the Gaussian Unitary, Orthogonal or Symplectic Ensembles (GUE,GOE, GSE). We are going to study time evolution and spread complexity of a generic state in the basis in which the Hamiltonian is drawn. Since the matrix is being drawn randomly, we can pick this to be the state $(1,0,0,\cdots)$ without loss of generality.
@@
$\frac{1}{4}V^{\prime}(\omega)=\mathrm{p.v.}\int dE\,\frac{\rho(E)}{\omega-E}\,$, (104)
@@
As explained in Balasubramanian:2022dnj, we can use (104) to calculate the density of states from the potential of the RMT, and then we can use (71) from Sec. 6.1.4 to determine the coarse-grained Lanczos spectrum and the spread complexity. This construction is averaged over the ensemble, but as shown in Balasubramanian:2022dnj, at large $N$ this gives the same result as coarse-graining over nearby Lanczos indices in a single draw from the ensemble, i.e., using the coarse-graining procedure outlined in Sec. 6.1.4. But we can do better than this and determine the distribution of Lanczos coefficients. For Gaussian RMTs, this distribution was worked out in Dumitriu_2002. The Jacobian arising from the basis transformation to get a tridiagonal Hamiltonian will be the same for any potential. The authors of Balasubramanian:2022dnj used this to show that for an RMT with an arbitrary potential like (102) the distribution of Lanczos coefficients is
@@
At large $N$ the distribution (105) is sharply peaked. So we can use the method of saddle points to extract the ensemble average (the saddle point) and the covariances (from the quadratic expansion around the saddle point). The ensemble average precisely reproduces (71), which we already derived through other means, and which applies to any theory for which we know the density of states. A formula for the covariances of the $a_{n}$ and $b_{n}$ is derived in Balasubramanian:2022dnj, Balasubramanian:2023kwd. We will not repeat the derivation here for lack of space, but the result is
$\mathrm{cov}(a_{i},a_{i+\delta})=4\mathrm{cov}(b_{i},b_{i+\delta})=\frac{1}{2\pi\beta_{D}N}\int_{0}^{2\pi}dk\,\frac{e^{2ik\delta}}{\lambda(k,a_{i},b_{i})}\,$, (106)
@@
$\lambda(k,a_{i},b_{i})=\int dE\frac{V^{\prime}(E)-V^{\prime}(a_{i})}{b_{i}(E-a_{i})}\eta((E-a_{i})/b_{i},e^{ik})\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \eta(x,t)=\frac{x}{\pi\sqrt{4-x^{2}}}\frac{1}{t+1/t-x}\,.$ (107)
@@
Among other things, these results show that the variance in $a(x)$ around the ensemble mean should be four times the variance in $b(x)$ for any RMT, where we are taking $x=n/N$ and coarse-graining as described in Sec. 6.1.4. The authors of Balasubramanian:2022dnj showed that these analytical formulae reproduce numerical results for RMTs with various potentials.
@@
Following Balasubramanian:2022tpr, the spread complexity of the time-evolved thermofield-double state for a GUE Hamiltonian is plotted in Fig. 9 for various matrix sizes $N$ and inverse temperatures $\beta$. The dynamics exhibit four characteristic regimes: a linear ramp up to a peak that is exponential in the entropy (and hence linear in $N$, the Hilbert space dimension), followed by a slope down to a plateau. The authors of Balasubramanian:2022tpr showed that the peak and slope arise from eigenvalue correlations known as spectral rigidity in chaotic theories. These correlations can be removed by drawing eigenvalues at random from the Wigner semicircle distribution – this produces a spectrum without eigenvalue repulsion, like the Poisson spectra expected for integrable theories. The evolution with such a spectrum (Fig. 9 in light hues) shows only the ramp and plateau. Thus, the peak and slope are characteristic of late time chaos. The same patterns appear in the late time spread complexity of the chaotic SYK model Balasubramanian:2022tpr.
@@
Figure 9: Quantum state complexity of the time-evolved thermofield-double state over an exponentially large period of time for different values of $N$ and $\beta$, as described in the main text. Dark Hues: GUE ensemble. Going from highest (blue) to lowest (yellow) curves we have $\beta=\{0,1,2,5,10\}$. In each case, we have plotted ensembles with $N=\{1024,1280,1536,1792,2048,2560,3072,3584,4096\}$. Complexity grows linearly to a peak, followed by a downward slope to a plateau. Light Hues: Ensemble with the same density of states as GUE, but without correlations between eigenvalues. In this case, the curves plateau without reaching a peak followed by a downward slope. Figure taken from Balasubramanian:2022tpr.
@@
Quantum chaotic systems are conjectured to have spectra that are well-described by RMT statistics Dyson_62_Statistical, PhysRevLett.52.1. In view of the results above, Balasubramanian:2023kwd conjectured that: Quantum chaotic systems display a Lanczos spectrum that is well described by RMT. As we discussed, the coarse-grained Lanczos spectrum can be determined from just the density of states. So the content of this conjecture lies in the correlations – it is saying that the covariances of the Lanczos coefficients are described by the RMT formulae stated above.
The conventional statements relating quantum chaos to Random Matrix Theory concern the statistics of nearby energy levels, and equivalently late time dynamics. Correspondingly, the universal RMT results that we described above concern the “bulk” of the Lanczos spectrum, i.e., for indices of $O(d)$ where $d$ is the dimension of the Hilbert space. In general Lanczos coefficients with larger indices control transitions to and between states that are further down in the Krylov chain that are populated later in the dynamics, although the precise connection between Lanczos index and time remains a topic of study. As such, the universal behavior we described applies to the late time dynamics. The Lanczos coefficients with indices that scale sub-extensively in $N$ must be computed separately by the recursion method and reflect non-universal and system-specific early time behavior even in systems that show late time chaos. On a related note, a precise setup where RMT emerges in holography is the SYK model, and the Thouless time in this setup was derived using the Spectral Form Factor (SFF) in Nosaka:2018iat (and is associated with the dip of SFF). Since the SFF plays a central role in spread complexity computations Balasubramanian:2022tpr, we also expect that the late time regime will correspond to the time after the peak of the spread complexity, and this was confirmed numerically in Balasubramanian:2022tpr, Balasubramanian:2023kwd.
To test this conjecture, we can numerically compute the Lanczos spectrum of integrable and chaotic spin chains such as
@@
where $X_{i}$, $Y_{i}$, $Z_{i}$ are Pauli spin operators on site $i$; $K$ is the number of spins; $A_{k}$, $B_{k}$ and $C_{k}$ are numerical parameters; and $H_{\text{site}}=0.5Z_{3}+0.3X_{3}$ is a one-site disorder operator that helps breaking parity and symmetry under the action of $\sum_{i}Z_{i}$. The model defined by $H_{\text{int1}}$ is integrable, while $H_{\text{cha1}}$ is chaotic as suggested by their level statistics, and have been explored in the literature Ba_uls_2011, Rabinovici:2022beu. Indeed, as shown in Balasubramanian:2023kwd, the analytical formulae developed above give an excellent description of the Lanczos coefficients in all cases, but the variances and covariances deviate significantly from the RMT formula for integrable spin chains, and in particular are much larger.
We can also write down an explicit formula for the height of the late time plateau in the spread complexity. First, let $\hat{n}=\sum_{n}n\lvert K_{n}\rangle\langle K_{n}\rvert$ be the position operator along the Krylov chain associated to an initial state $\lvert\phi\rangle$. Then, define the complexity of an energy eigenstate $\lvert E\rangle$ relative to $\lvert\phi\rangle$ as $\mathcal{C}_{E}=\langle E\rvert\hat{n}\lvert E\rangle$. A short calculation in Balasubramanian:2023kwd shows that the height of the late time plateau in the spread complexity will be
@@
where $p(E)=|\langle E|\phi\rangle|^{2}$ is the initial probability of eigenstate $\lvert E\rangle$. If $p(E)$ is continuous, we can replace $\mathcal{C}_{E}$ by the local average in a small window over energies to get a coarse-grained eigenstate complexity $\overline{\mathcal{C}_{E}}$. A calculation in Balasubramanian:2023kwd shows that $\overline{\mathcal{C}_{E}}$ is entirely determined by Lanczos coefficients, which are in turn determined in terms of the density of states following Eq. (71). In terms of the coarse grained eigenstate complexities, and in the large system limit, we can take a continuum limit of (109), which yields
$\mathcal{C}_{\rm plateau}=\int dE\,\rho(E)\,p(E)\,\overline{\mathcal{C}_{E}}\,.$ (110)
This means that for large systems and initial states with continuous support in the energy basis, the plateau of the spread complexity will not discriminate between chaotic and integrable theories with the same average density of states. However, if $p(E)$ is “noisy” rather than smooth, and the noise is negatively correlated with the noise in $\mathcal{C}_{E}$, the plateau can be significantly lower than the averaged value in (110). This lowering can be larger for integrable theories because, as we discussed above, the variance in the Lanczos is higher, leading to a form of Anderson localization of energy eigenstates along the Krylov chain (see also Rabinovici:2022beu). So if an eigenstate is “centered” further from the beginning of the chain it tends to have less overlap with the initial state for an integrable theory. It would be interesting to investigate further how tuning the initial state leads to differences in the late time behavior for integrable versus chaotic theories because of variances and covariances in the Lanczos spectrum
These analyses have been extended to a broad class of integrable and chaotic models, including the SYK model and spin chains. Key aspects such as the presence or absence of the peak and slope Erdmenger:2023wjg, the peak height Baggioli:2024wbz, and the system’s approach to the plateau Alishahiha:2024vbf have been advocated as indicators of chaos. However, it seems that Krylov complexity Bhattacharjee:2022vlt and spread complexity Huh:2023jxt are not sensitive to saddle-dominated scrambling (exponential growth due to unstable saddle points in phase space), at least for the initial states/operators studied in these papers, a phenomenon that OTOCs also fail to distinguish from genuine quantum chaos Xu:2019lhc.
@@ 6.4 Open problems and directions
These quantities are known exactly in some QFTs. For example, in 2D CFTs, thermal two-point correlators are universally determined by Ward identities and are the same for both integrable and chaotic CFTs. In this case, there is an infinite sequence of $b_{n}$ Lanczos coefficients that grow linearly for large $n$ dymarsky2021krylov. At least at first sight, this serves as a counterexample to the universal operator growth hypothesis. This issue was explored in Avdoshkin:2022xuw, Camargo:2022rnt, where the authors argued that to meaningfully characterize operator growth in QFTs, one must first introduce appropriate cut-offs. They analyzed the effects of both UV and IR cut-offs, and concluded that the chaotic or scrambling properties of QFTs may be captured by $b_{n}$’s larger than the UV cut-off $\Lambda$. Meanwhile, the IR cut-off (e.g. modeled by introducing a mass) induced a separation (staggering) between even and odd Lanczos coefficients that grew linearly with different intercepts Camargo:2022rnt.
Thus, distinguishing characteristic features of chaotic and integrable operator growth in continuum QFTs remains an open problem. A promising approach could be to employ von Neumann algebras and algebraic QFTs. Some progress in this direction has been made in Ouseph:2023juq, Gesteau:2023rrx, but a direct connection with Krylov-basis methods has not been established. Alternatively, we could compute the covariances of the Lanczos coefficients in QFT, in view of the conjecture in Balasubramanian:2022dnj, Balasubramanian:2023kwd that their structure separates chaos from integrability. In theories with a discrete spectrum, it may also help to project onto a microcanonical band of energies (see also Kar:2021nbm) thereby rendering the accessible Hilbert space finite dimensional, thus enabling us to directly adopt the ideas and methods that have been developed for the latter.
@@
The relation between Krylov and Nielsen’s approaches has been debated in several works. In Lv:2023jbv, the authors argued that Krylov complexity can be interpreted geometrically as a distance in a certain projection of the Fubini-Study geometry that is natural in quantum optics, making it amenable to a circuit complexity interpretation. However, as we already saw from the symmetry-based examples, Krylov complexity, as a function of time, is proportional to the volume in the Fubini-Study geometry. As a volume, rather than a length, it does not respect some properties of geodesics, such as the triangle inequality. Such properties were suggested in Aguilar-Gutierrez:2023nyk as an obstruction to Krylov complexity serving as a distance measure. However, it is not clear that the Fubini-Study setting is the one in which we should seek a relation with Nielsen’s complexity. Indeed, Craps:2023ivc argued that, in certain setups, there is a relation between the time average of spread complexity of state evolution and an upper bound on Nielsen’s complexity with specific fine-tuned penalty factors. Recalling that the Nielsen approach has ambiguities in the choice of basis of generators and cost function, it remains possible that some judicious choice could establish a direct relation between the Nielsen and Krylov/spread approaches to complexity, at least for some specific choices of states. Perhaps useful progress could be achieved by formulating Krylov/spread complexity approach to general, time-dependent Hamiltonians (see progress in Takahashi:2024hex). This would be closer in spirit to the starting point in Nielsen’s definitions and could allow us to use bounds on circuit complexity to constrain Krylov complexity. Indeed, if we treat the evolution parameter as a circuit time, we can mathematically think about the complexity of formation of states in the language of spread complexity Caputa:2022eye. Thus it remains an interesting problem to systematically relate the complexity of state/operator spreading to circuit complexity.
@@
It would be interesting to extend the spread/Krylov complexity approaches to new systems. First, the generalization to time-dependent Hamiltonians is important to achieve for the reasons described above, and because we would like to understand systems that are driven by external sources. Second, if we contemplate external sources, we should also consider open quantum systems, whose dynamics are controlled by Lindbladian evolution of the density matrix Carolan:2024wov. There is a substantial literature on the Linbladian formalism which is partly reviewed in Nandy:2024htc. The interaction of an external bath with an open system certainly affects the structure of entanglement spreading (see, e.g., Agon:2014uxa), and has generally been a subject of significant recent interest in condensed matter physics through the study of measurement-driven phase transitions (see, e.g., block2022measurement). The spread complexity of measurement-induced non-unitary dynamics has been explored in Bhattacharya:2023yec. Meanwhile, Das:2024zuu have uncovered various bounds and relations pertaining to the state/operator complexity of purifying a mixed state. Another interesting avenue is to study the spread complexity of Carrollian theories of relevance for flat space holography. Likewise it would be interesting to study the kind of chiral theories that appear in the AdS/CFT dual to highly rotating extremal black holes Guica:2008mu, Balasubramanian:2003kq, Balasubramanian:2009bg and the matrix models that appear as the dual to the 11-dimensional M-theory Banks:1996vh, Balasubramanian:1997kd, Polchinski:1999br. In this direction, spread complexity has been studied in the planar limit of holographic theories in Das:2024tnw.
@@
In Sec. 7.5 we will discuss how spread complexity of the double scaled SYK model is directly related to the length of the dual wormholes in two-dimensional Jackiw-Teitelboim gravity. But we could ask a broader question. Many lines of argument suggest that the microstates of black holes are immensely complex and may appear to be hidden behind a horizon simply because simple probes cannot discriminate between them – see, e.g., Balasubramanian:2005mg, Balasubramanian:2005kk, Balasubramanian:2006jt, Balasubramanian:2006iw. Similarly, a number of authors have argued that in the AdS/CFT correspondence excitations behind the horizon are encoded in the CFT by exponentially complex operators Engelhardt:2021mue, engelhardt2022finding in a non-isometric manner Akers:2022qdl. The information stored in Hawking radiation is also supposed to be encoded in an extremely complex manner, with the complexity expressed in the dual gravitational theory in terms of geometric obstructions preventing asymptotic observers from accessing the information Brown:2019rox. At least in toy models of black hole evaporation like Penington:2019kki, such considerations can be explicitly realized and characterized in information-theoretic and complexity-theoretic terms Balasubramanian:2023xdp. But black holes can form from collapse of simple states like an ordered cloud of particles. So, the immense complexity of black hole microstates, and of the subsequent encoding of the initial state in the Hawking radiation, must be generated dynamically by evolution under the Hamiltonian of the theory. An important question is to understand how this happens. Perhaps the techniques of Krylov/spread complexity, or of Nielsen complexity from the previous section, can help. One recent result shows that if a holographic CFT time evolution operator is approximately pseudorandom or Haar random on a low energy subspace of states, there must be an event horizon in the dual theory of gravity Engelhardt:2024hpe. Likewise, it seems that any region of AdS with sufficiently high holographic encoding complexity will be robustly protect from low-rank measurements, and hence be inaccessible to low-complexity observers Balasubramanian:2023xdp. It would be interesting to understand how the dynamical evolution of simple states can produce complex states with such properties.